OpenStudy (ninabina):

The perimeter of the rectangle is 24 and the perimeter of the triangle is 23. The equations for the triangle is x+1, x+y, and y+6. The equations for the rectangle is x and 3y. Write an equations to that represents the perimeter of the triangle. Write an equation in standard form the represents the perimeter of the rectangle. Write a system of linear equations to represent the problem. The solve the system of linear equations to find x and y.

OpenStudy (anonymous):

alright, I can help you out with the first part, then you'll be able to figure out the last pasrt, okay? ^_^ so: Given: P(rectangle) = 24 L = 3y w = x General perimeter equation of rectangle = 2L + 2w P = 2(3y) +2x = 6y+2x -------------------------- P(triangle) = 23 S1(side1) = x+1 S2(side2) = x+y S3(side3) = y+6 General equation of triangle's perimeter = S1 + S2 + S3 P = x+1 + x+y + y+6 = 2x +2y + 7 Now if you understand what I have done here, then you'll be able to proceed with the second part ^_^ good luck~ :)